0.3x-0.2y=4
0.5x+0.3y=1

Answer by mananth(12270)

0.3x-0.2y=4
0.5x+0.3y=1
multiply both equations by 10
we get
3 x -2 y = 40 .............1
5 x + 3 y = 10 .............2
Eliminate y
multiply (1)by 3
Multiply (2) by 2
9 x -6 y = 120
10 x + 6 y = 20
Add the two equations
19 x = 140
/ 19
x = 7.37
plug value of x in (1)
3 x + -2 y = 40
22.11 + -2 y = 40
-2 y = 40-22.11
-2 y = 17.89
y = -8.95